OpenStudy (anonymous):

Find the quotient. 48a^3bc^2 ÷ 3abc

OpenStudy (anonymous):

Add exponents that have the same base when multiplying, subtract when dividing: \[\frac{x^3}{x^2}=x^{3-2}=x\] \[16 a^4 b^2 c^3\]

OpenStudy (anonymous):

However, I think you also might have a typo you may want to check... If you meant the below, it's not the same answer. Got to watch parenthesis.\[\frac{(48 a^3 (b c^2))}{(3 a b c)} = 16 a^2 c\]

OpenStudy (unklerhaukus):

\[ 48a^3bc^2 ÷ 3abc=\frac{ 48a^3bc^2}{3abc}\]\[=\frac {48}{16}a^{3-1}b^{1-1}c^{c-1}\]

OpenStudy (anonymous):

there were no () in the prob

OpenStudy (anonymous):

If there were none, then you'll be using the first answer I posted. Double check though just in case. It's a common mistake. Btw, just to be clear: \[\frac{(48a^3(bc^2))}{(3abc)}=\frac{48a^3bc^2}{3abc}\]

OpenStudy (anonymous):

You can type a fraction like that by using \frac {numerator}{denominator} \[ \frac{numerator}{denominator} \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

whats the quotient of \48a 3 bc 2 ÷ 3abc